# Style Ambiguity Loss Without Labeled Datasetss

### v2

@@ 5.1.1 Comparison with GAN Baseline
    Image Dimensionality: 256 or 512
@@
    Dataset: Full or Mediums
@@
    Batch Size: 256 or 128
@@
    Gradient Penalty: GP (We do use GP) or No GP (We do not use GP)
This gave us sixteen combinations, and so we had 16 generators. Scores are shown in table 3. Refer to appendix A for box plots visualizing these results.
Model AVA IR
Image Dim 256; Full Dataset
Batch 256 No GP 3.53 ( 0.13 ) -2.23 ( 0.02 )
Batch 256 GP 3.73 ( 0.17 ) -2.14 ( 0.03 )
Batch 128 No GP 3.9 ( 0.16 ) -1.39 ( 0.23 )
Batch 128 GP 3.79 ( 0.16 ) -1.29 ( 0.23 )
Image Dim 512; Full Dataset
Batch 256 No GP 4.1 ( 0.16 ) -2.13 ( 0.05 )
Batch 256 GP 3.64 ( 0.08 ) -2.2 ( 0.02 )
Batch 128 No GP 4.25 ( 0.08 ) -1.78 ( 0.1 )
Batch 128 GP 4.21 ( 0.14 ) -0.89 ( 0.1 )
Image Dim 256; Mediums Dataset
Batch 256 No GP 3.88 ( 0.19 ) -1.46 ( 0.18 )
Batch 256 GP 3.85 ( 0.13 ) -1.6 ( 0.18 )
Batch 128 No GP 4.38 ( 0.23 ) -1.64 ( 0.18 )
Batch 128 GP 4.04 ( 0.14 ) -1.39 ( 0.09 )
Image Dim 512; Mediums Dataset
Batch 256 No GP 4.31 ( 0.24 ) -1.53 ( 0.12 )
Batch 256 GP 3.97 ( 0.14 ) -1.31 ( 0.12 )
Batch 128 No GP 4.34 ( 0.2 ) -1.22 ( 0.13 )
Batch 128 GP 4.23 ( 0.12 ) -1.2 ( 0.14 )
Table 3: Caption
The GANs did not perform well. The highest AVA score attained by a GAN was 4.34, which was lower than the AVA score for all of the diffusion models. The highest IR score was -0.89, which was higher than the IR score for disc-med, clip-full and disc-full, but the average GAN score was still far below that of the diffusion models. Evidently, GANs do not seem to perform very well. Their inferiority is compounded by the problems of training GANs, such as mode collapse and instability (Arjovsky & Bottou, 2017). The much lower standard deviation of aesthetic scores for the GANs also strongly indicates mode collapse.
5.2 User Study
We also conducted a user study to evaluate a subset of the models. The goal of this study was to evaluate each model for human appreciation, as we had done with IR and AVA, as well as novelty, surprise, ambiguity and complexity. For each model we did evaluate in this user study, we generated five images with the same seeds. Then, for each image, users to gauge 5 metrics, by asking them each a question, as was done in Elgammal et al. (2017). These 5 metrics were based on the components of aesthetic arousal (Berlyne, 1971), that being the level of alertness or excitation that an image provokes in a viewer. "Too much" arousal can actually have a negative effect on subjective appreciation. The metrics and their associated questions are as follows:
  1. 1.
    Appreciation: How much do you like this image: 1-extremely dislike, 2-dislike, 3-Neutral, 4-like, 5-extremely like.
  2. 2.
    Novelty: Rate the novelty of the image: 1-extremely not novel, 2-some how not novel, 3-neutral, 4 somehow novel, 5-extremely novel.
  3. 3.
    Surprise: Do you find the image surprising: 1-extremely not surprising, 2-some how not surprising, 3-neutral, 4-some how surprising, 5-extremely surprising.
  4. 4.
    Ambiguity: Rate the ambiguity of the image. I find this image: 1-extremely not ambiguous, 2-some how not ambiguous, 3-neutral, 4-some how ambiguous, 5-extremely ambiguous.
@@ 5.2 User Study
    Complexity: Rate the complexity of the image. I find this image: 1-extremely simple, 2-some how simple, 3-neutral, 4-somehow complex, 5-extremely complex
@@
Table 4: User Study Results
Results are shown in table 4. We see here that kmeans-med outperforms all of the other models on all of the metrics. For both datasets, the lowest Appreciation score was attained by the baseline discriminator model. For the novelty metric, which is a proxy for the "creative" aspect of creativity, both discriminator-free methods outperform the discriminator using the mediums datasets. For the other metrics, which are less important but still interesting given that they are components of aesthetic arousal, the discriminator methods still underperform. Complexity for example, shows that for both datasets, the discriminator performs the worst.
5.3 Visual Results
We also provide a few visual results in figure 1. Each image in each row was generated with the same prompt and random seed. The leftmost column has the prompt used to generate each picture in each row, and each column is labeled with the model used to generate the images in said column. For more pictures, consult appendix B.
@@ 5.3 Visual Results
Both discriminator model, as well as clip-full generate blurry, abstract patterns. While those are certainly interesting, they do not match the prompt. Kmeans-full also suffers from a proclivity to draw mens’ faces, regardless of prompt. However, both clip-med and kmeans-med demonstrate prompt alignment and interesting visuals.
5.4 Content and Style Similarities
We also compared the similarities between each model. Given that the nth image generated by each model used the same prompt and seed across all models, we could compare the cosine distance of the embeddings of the nth image for each model with the embeddings of the nth image of every other model, and average them. We used style and content embeddings from the dino-vits16 checkpoint (Caron et al., 2021) from https://huggingface.co/facebook/dino-vits16. Many other works have commented on the ability to extract separable style and content from vision transformers (Tumanyan et al., 2022; Kwon & Ye, 2023).
disc-full clip-full kmeans-full disc-med clip-med kmeans-med
disc-full 1.0 0.29 0.26 0.29 0.27 0.29
clip-full 0.29 1.0 0.24 0.41 0.27 0.26
kmeans-full 0.26 0.24 1.0 0.24 0.38 0.31
disc-med 0.29 0.41 0.24 1.0 0.25 0.26
clip-med 0.27 0.27 0.38 0.25 1.0 0.33
kmeans-med 0.29 0.26 0.31 0.26 0.33 1.0
Table 5: Average Content Similarities
disc-full clip-full kmeans-full disc-med clip-med kmeans-med
disc-full 1.0 0.23 0.23 0.22 0.27 0.27
clip-full 0.23 1.0 0.28 0.45 0.32 0.3
kmeans-full 0.23 0.28 1.0 0.21 0.59 0.4
disc-med 0.22 0.45 0.21 1.0 0.26 0.25
clip-med 0.27 0.32 0.59 0.26 1.0 0.49
kmeans-med 0.27 0.3 0.4 0.25 0.49 1.0
Table 6: Average Style Similarities
5.5 Visualizing The Possibility Space
For each model, we combined each pair of evaluation datasets and then then performed dimensionality reduction using PCA and TSNE to embed images into 2 dimensions. Thus, we could visualize the overlap, or lack thereof, between the possibility space of each model.
Figure 2: Possibility Space
This visualization, as well as the similarity tables, proves that all the models being studied are creating images in very distinct spaces.
6 Conclusion
@@
In this paper, we introduce versatile CLIP-based and K-Means-based creative style ambiguity losses, which eliminate the need for training a separate style classifier on labeled datasets. We have demonstrated that these new style ambiguity losses are just as good, oftentimes better, than the baseline discriminator based style ambiguity loss, as shown by automated metrics and user studies. Furthermore, this method can likely be applied to different datasets, consisting of different images or even different modalities like audio or video. Using the K-Means Classifier for style ambiguity loss is particularly adept at this, given there is no need for any labels. However, K-Means is just one unsupervised clustering method, and others may be worth exploring as well.
Broader Impact Statement
Many are concerned about the impacts of generative AI. By making art, this work infringes upon a domain once exclusive to humans. Companies have faced scrutiny for possibly using AI (Gutierrez, 2024), and many creatives, such as screenwriters and actors, have voiced concerns about whether their jobs are safe (del Barco, 2023). Nonetheless, using AI can help humans by making them more efficient, providing inspiration, and generating ideas (Fortino, 2023; Campitiello, 2023; Darling, 2022). It’s also not certain how copyright protection will function for AI-generated art (Watiktinnakorn et al., 2023), given copyright law is based on the premise that creative works originate solely from human authorship. Clear, consistent policies, both at the government level and by industry and/or academic groups, will be needed to mitigate the harm and maximize the benefits for all members of society.
7 Assistance
Author Contributions
This work was done without any outside assistance or collaboration.
Acknowledgements
Rutgers Office of Advanced Research Computing kindly provided the computing infrastructure to run the experiments. A special thanks goes out to Dr. Ahmed Elgammal and Dr. Eugene White for their past advice prior to this work.
Appendix A GAN Results
Box plots for the image reward and aesthetic scores are shown in tables 7 and 8, respectively. Some example images are shown in figure 3
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
Table 7: Image Reward
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
Table 8: Aesthetic Reward
Figure 3: GAN Examples
Appendix B Additional Images
We display a few more images in figures 4 and 5. Once again, each image in each row was generated with the same seed and same prompt.
Figure 4: More Images
Figure 5: More Images
Appendix C WikiArt Classes
We host the original WikiArt-Full dataset on REDACTEDWHILEUNDERBLINDREVIEW. The 27 WikiArt style classes, $L_{full}$ are listed in table 9
contemporary-realism art-nouveau-modern abstract-expressionism
northern-renaissance mannerism-late-renaissance early-renaissance
realism action-painting color-field-painting
pop-art new-realism pointillism
expressionism analytical-cubism symbolism
fauvism minimalism cubism
romanticism ukiyo-e high-renaissance
synthetic-cubism baroque post-impressionism
impressionism rococo na-ve-art-primitivism
Table 9: Styles
Appendix D Medium Subset
D.1 Captions
All captions were generated using the Salesforce/blip-image-captioning-base checkpoint downloaded from https://huggingface.co/Salesforce/blip-image-captioning-base.
D.2 Subset
The categories of images that had the highest percentage of images in the categories with text captions that contained one of the words in the relevant text prompt set, as well as the percentage of images that did so, and the quantity of images in said category are shown in table 10 . We host the Mediums datasets at REDACTEDWHILEUNDERBLINDREVIEW.
style class quantity percentage
expressionism 6054 91.56
post-impressionism 5832 89.25
fauvism 841 96.08
abstract-expressionism 2518 89.95
na-ve-art-primitivism 2148 93.39
cubism 2027 91.61
synthetic-cubism 197 89.85
analytical-cubism 105 91.43
new-realism 280 96.07
action-painting 93 93.55
Table 10: Mediums
Appendix E Training
For reproducibility and transparency, the hyperparameters are listed in table 11 and table 12. All experiments were implemented in Python, building the models in pytorch (Paszke et al., 2017) using accelerate (Gugger et al., 2022) for efficient training. The diffusion models also relied on the trl (von Werra et al., 2020), diffusers (von Platen et al., 2022) and peft (Mangrulkar et al., 2022) libraries. The K-Means clustering was done using the k means implementation from scikit-learn (Pedregosa et al., 2011). A repository containing all code can be found on github at REDACTEDWHILEUNDERREVIEW. Each experiment was run using two NVIDIA A100 GPUs with 40 GB RAM.
Hyperparameter Value
Epochs 25
Effective Batch Size 8
Batches per Epoch 32
Inference Steps per Image 30
LORA Matrix Rank 4
LORA $\alpha$ 4
Optimizer AdamW
Learning Rate 0.0015
AdamW $\beta_{1}$ 0.9
AdamW $\beta_{2}$ 0.99
AdamW Weight decay 1e-4
AdamW $\epsilon$ 1e-8
Table 11: DDPO Hyperparameters
Hyperparameter Value
Epochs 100
Batch Size 32
Optimizer Adam
Learning Rate 0.001
Adam $\beta_{1}$ 0.9
Adam $\beta_{2}$ 0.99
Adam Weight decay 0.0
Adam $\epsilon$ 1e-8
Noise Dim 100
Wasserstein $\lambda$ 10
Leaky ReLU negative slope 0.2
Convolutional Kernel 4
Convolutional Stride 2
Transpose Convolutional Kernel 4
Transpose Convolutional Stride 2
Table 12: CAN Hyperparameters
E.1 Architecture
For diffusion model training, the text encoder, autoencoder and unet were all loaded from https://huggingface.co/stabilityai/stable-diffusion-2-base. These model components were all frozen, but we added trainable LoRA weights to the cross-attention layers of the Unet. Parameter counts are shown in table 13. The diffusion model components used the same amount of parameters regardless of image size, but the generator and discriminator had more parameters as image size increased.
Model Component Total Parameters Trainable Parameters Percent Trainable
Text Encoder 34,0387,840 0 0%
Autoencoder 83,653,863 0 0%
UNet 866,740,676 829,952 0.1%
Generator (Image Dim 512) 48,014,784 48,014,784 100%
Discriminator (Image Dim 512) 20,115,932 20,115,932 100%
Table 13: Parameter Counts
We used the convolutional neural network (Dumoulin & Visin, 2018) architecture described in Elgammal et al. (2017) for the CAN but had to use more/less layers to produce higher/lower dimension images. The generator takes a $1\times 100$ gaussian noise vector $\in\mathbb{R}^{100}\sim\mathcal{N}(0,I)$ and maps it to a $4\times 4\times 2048$ latent space, via a convolutional transpose layer with kernel size = 4 and stride =1, followed by 6 transpose convolutional layers each upscaling the height and width dimensions by two, and halving the channel dimension (for example one of these transpose convolutional layers would map $\mathbb{R}^{4\times 4\times 2048}\rightarrow\mathbb{R}^{8\times 8\times 1024}$) followed by batch normalization (Ioffe & Szegedy, 2015) and Leaky ReLU (Maas et al., 2013), and then one final convolutional transpose layer with output channels = 3 and tanh (Dubey et al., 2022) activation function. Diagrams of the generator is shown in the figure 6.
For the discriminator, we first applied a convolution layer to downscale the input image height width dimensions by 2 and mapped the 3 input channel dimensions to 32 ($\mathbb{R}^{512\times 512\times 3}\rightarrow\mathbb{R}^{256\times 256\times 32}$) with Leaky ReLU activation. Then we had 5 convolutional layers each downscaling the height and width dimensions by 2 and doubling the channel dimension (for example, one of these convolutional layers would map $\mathbb{R}^{256\times 256\times 32}\rightarrow\mathbb{R}^{128\times 128\times 64}$) with batch normalization and Leaky ReLU activation. Then we had two more convolutional layers, each downscaling the height and width dimensions but keeping the channel dimensions constant (using the prior layer’s channel dimensions), with batch normalization and Leaky ReLU activation. The output of the convolutional layers was then flattened. The discriminator had two heads- one for style classification (determining which style a real image belongs to) and one for binary classification (determining whether an image was real or fake). The binary classification head consisted of one linear layer with one output neuron. The style classification layer consisted of 2 linear layers with LeakyReLU activation and Dropout, with output 1024 output neurons and 512 output neurons, respectively, followed by a linear layer with 27 output neurons for the 27 artistic style classes. A diagram of the discriminator is shown in 7.
Figure 6: Generator Architecture (Image Dim 512)
Figure 7: Discriminator Architecture (Image Dim 512)

### v3

@@ IV-B User Study
    Appreciation (App): How much do you like this image: 1-extremely dislike, 2-dislike, 3-Neutral, 4-like, 5-extremely like.
@@
    Novelty (Nov): Rate the novelty of the image: 1-extremely not novel, 2-some how not novel, 3-neutral, 4 somehow novel, 5-extremely novel.
@@
    Surprise (Sur): Do you find the image surprising: 1-extremely not surprising, 2-some how not surprising, 3-neutral, 4-some how surprising, 5-extremely surprising.
@@
    Ambiguity (Amb): Rate the ambiguity of the image. I find this image: 1-extremely not ambiguous, 2-some how not ambiguous, 3-neutral, 4-some how ambiguous, 5-extremely ambiguous.
@@
    Complexity (Com): Rate the complexity of the image. I find this image: 1-extremely simple, 2-some how simple, 3-neutral, 4-somehow complex, 5-extremely complex
@@
TABLE II: User Study Results
Results are shown in table II. We see here that kmeans-med outperforms all of the other models on all of the metrics. For both datasets, the lowest Appreciation score was attained by the baseline discriminator model. For the novelty metric, which is a proxy for the ”creative” aspect of creativity, both discriminator-free methods outperform the discriminator using the mediums datasets.
IV-C Visual Results
We also provide a few visual results in figure 1. Each image in each row was generated with the same prompt and random seed. The leftmost column has the prompt used to generate each picture in each row, and each column is labeled with the model used to generate the images in said column.
Both discriminator model, as well as text-full generate blurry, abstract patterns. While those are certainly interesting, they do not match the prompt. However, both text-med and kmeans-med demonstrate prompt alignment and interesting visuals.
@@ IV-C Visual Results
V Conclusion
@@
In this paper, we introduce versatile CLIP-based and K-Means-based creative style ambiguity losses, which eliminate the need for training a separate style classifier on labeled datasets. We have demonstrated that these new style ambiguity losses are just as good, oftentimes better, than the baseline discriminator based style ambiguity loss, as shown by automated metrics and user studies. This method can be applied to different image datasets or even different modalities like audio or video. Using the K-Means Classifier, or any unsupervised clustering method, for style ambiguity loss is particularly adept at this, given there is no need for any labels.